# Embedding Ontologies via Incorporating Extensional and Intensional Knowledge

### v4

@@ Abstract
Ontologies contain rich knowledge within domain, which can be divided into two categories, namely extensional knowledge and intensional knowledge. Extensional knowledge provides information about the concrete instances that belong to specific concepts in the ontology, while intensional knowledge details inherent properties, characteristics, and semantic associations among concepts. However, existing ontology embedding approaches fail to take both extensional knowledge and intensional knowledge into fine consideration simultaneously. In this paper, we propose a novel ontology embedding approach named EIKE (Extensional and Intensional Knowledge Embedding) by representing ontologies in two spaces, called extensional space and intensional space. EIKE presents a unified framework for embedding instances, concepts and their relations in an ontology, applying a geometry-based method to model extensional knowledge and a pretrained language model to model intensional knowledge, which can capture both structure information and textual information. Experimental results show that EIKE significantly outperforms state-of-the-art methods in three datasets for both triple classification and link prediction, indicating that EIKE provides a more comprehensive and representative perspective of the domain.  [^1] This paper is accepted by Data Intelligence.
keywords:
Ontology; Ontology embedding; Extensional knowledge; Intensional knowledge; Pre-trained language model
journal: Data Intelligence
@@ 1 Introduction
Ontologies, such as YAGO [33], WordNet [21] and BioPortal [27], serve as pivotal tools for organizing, representing, and semantically linking heterogeneous information. They provide a structured framework that defines concepts and their relationships, offering a structured knowledge representation that supports various applications like information retrieval [22, 16] and semantic web technologies [15, 6].
@@
Within an ontology, knowledge can be divided into two categories, intensional knowledge, acting as a terminology or taxonomy, describing general properties, characteristics of concepts and their semantic relationships, and extensional knowledge, also known as assertional knowledge, which pertains to specific individuals within the domain and provides type information about the concrete instances that belong to a specific concept [23, 6]. Figure 1 shows a sketch of such an ontology. Comprehensively representing both extensional knowledge and intensional knowledge within an ontology is essential for a nuanced, accurate and adaptable understanding of the domain [6].
Ontology embedding involves transforming each component in an ontology into continuous and low-dimensional vectors while preserving their semantic relationships. Recently, ontology embedding techniques have been widely investigated and applied to various applications such as semantic search and reasoning [1]. Integrating both extensional knowledge and intensional knowledge enriches ontology embedding and provides a more comprehensive and representative perspective of the domain, resulting in embeddings that better capture semantic relationships and similarities among instances and concepts, as well as better apply to applications such as materialization and classification. However, existing ontology embedding models fail to take both of these two types of knowledge into fine consideration simultaneously. For example, existing geometry-based approaches such as [20, 43] represent extensional knowledge well, but perform poorly in capturing intensional knowledge since they ignore the inherent characteristics and properties of the concepts. Especially, they can’t capture lexical and textual information. Concept2Vec [28] utilizes a random walk strategy to learn concept embedding through the similarity between concepts, ignoring the influence of a large number of instances. OWL2Vec* [5] considers the semantic aspects of graph structures, lexical information, and logical constructors, making it suitable for modeling intensional knowledge within an ontology. However, despite the fact that OWL2Vec* takes instances into account, its random walk strategy fail to clearly distinguish the differences between concepts and instances, making it hard to sufficiently model extensional knowledge.
@@ 2.1 Instance-level ontology embedding
Early knowledge embedding methods, called factual knowledge graph embedding (KGE), treat both concepts and instances as instances. Mainstream models are mainly divided into translational distance models such as TransE [2], TransH [38], TransR [19], TransD [14], and bilinear models, such as RESCAL [25], HolE [24], DistMult [40] and ComplEx [34]. Recently, more complicated KGE methods are proposed such as using box embeddings [30] and gaussian embeddings [12]. [36] provides a comprehensive review of these KGE models.
Some recent works apply pre-trained language models (PLMs) to knowledge embedding to capture textual information. For example, PretrainKGE [45] utilizes BERT [7] as encoders to derive embeddings of the entities and relations with descriptions, and input them into structured-based knowledge graph embedding models for training. KEPLER [37] uses PLMs as encoders to derive description-based embeddings and is trained on the objectives of knowledge embedding (KE) and masked language modeling (MLM). [39, 4] both apply sequence-to-sequence framework and pretrained language models to Knowledge Graph Completion (KGC). More recently, [42] explores Large Language Models for KGC. These methods can leverage descriptive information in knowledge graphs but cannot sufficiently represent concepts and model hierarchical structure.
@@ 2.2 Concept-level ontology embedding
Concept-level ontology embedding methods consider both instances and concepts in an ontological knowledge base as concepts. For instance, EL Embedding [18] constructs specific scoring functions and loss functions from logical axioms from EL++ by transforming logical relations into geometric relations. Onto2Vec [31] and OPA2Vec [32] are two ontology embedding algorithms based on word embeddings using skip-gram or CBOW architectures. Concept2Vec [28] utilizes a random walk strategy to learn concept embeddings through the similarity between concepts. CosE [9] utilizes angle-based and distance-based approaches to design two scoring functions to learn SubClassOf and DisjointWith while preserving transitivity and symmetry. Some works study embedding taxonomies and hierarchical structures. For example, [8] proposes n-ball to embed the taxonomy, and there exist a series of researches that utilize non-euclidean spaces, e.g., Poincaré embeddings [26]. Concept-level ontology embedding focuses on ontological concepts, disregarding instances within the knowledge base, hence its limited scope. Although some ontology embedding methods like OWL2Vec* [5] take instances into account, they fail to distinguish concepts and instances clearly because of the nature of their learning strategies like random walk. In contrast, our primary focus revolves around methodologies that differentiate concepts and instances.
@@ 2.3 Differentiating concepts and instances for ontology embedding
Recently, some studies have attempted to differentiate between concepts and instances in knowledge representation learning. TransC [20] represents each concept as a spherical region and each instance as a point vector in the same vector space, capable of handling the isA transitivity. TransFG [44] adopts TransC’s geometric modeling approach but embeds instances and concepts into different vector spaces to address instance ambiguity. To address the anisotropy issue in TransC and TransFG, [43] introduces TransEllipsoid and TransCuboid, representing concepts respectively as ellipsoidal and cuboidal regions. Additionally, JOIE [11] and DGS [13] learn embeddings from two-view knowledge graphs, yet they lacks interpretability for the isA transitivity. [46] and RMPI [10] show utilizing ontology-view information can enhance instance-view knowledge representation for applications like knowledge graph completion and inductive relation inference.
Different from previous methods like TransFG [44] and JOIE [11] which embed ontologies in two spaces, one space only for concepts and the other only for instances, EIKE embeds extensional knowledge and intensional knowledge in two distinct spaces respectively. This approach represents ontologies more sufficiently since it combine structure-based embeddings and textual-based embeddings. Additionally, EIKE is the first to finely consider two aspects of semantics of a concept: (1) a hight level-summary of its instances; (2) its inherent properties, characteristics and associations with other concepts.
@@ 3 Task Definition
Ontology. Due to differences in logical constraints and conceptual relations, ontologies have varying expressivity, such as DL-Lite [3] and $\mathcal{EL}$ [17]. In this work, following [20, 43], we focus on a simple and widely used ontology where there is only one kind of relation, i.e., SubClassOf, between atomic concepts. Ontology is simply formalized as $O=\{C,I,R,T,L\}$. $C$ and $I$ denote the sets of concepts and instances respectively. Relation set $R$ is categorized into three types as $R=\{r_{e},r_{c}\}\cup R_{l}$, where $r_{e}$ and $r_{c}$ denote InstanceOf relation and SubClassOf relation respectively, and $R_{l}$ is the relation set containing instance relations. The triple set $T$ is divided into three disjoint subsets: Relational triple set $T_{l}=\{(i_{m},r_{l},i_{n})\mid i_{m},i_{n}\in I,r_{l}\in R_{l}\}$, SubClassOf triple set $T_{c}=\{(c_{m},r_{c},c_{n})\mid c_{m},c_{n}\in C\}$ and InstanceOf triple set $T_{e}=\{(i,r_{e},c)\mid i\in I,c\in C\}$. $L$ denotes the set of existing descriptions about certain concepts in ontology.
@@
Task Definition. We embed ontologies in extensional space and intensional space. Given an ontology $O$, we map each instance $i$ into a point vector $\mathbf{i}^{e}\in\mathbb{R}^{d}$ and each concept c into a geometric region ${G}(\mathbf{c},\mathbf{b})$ in extensional space and map $\mathbf{i}^{e}$ into a point vector $\mathbf{i}^{i}\in\mathbb{R}^{d}$ and concept $c$ into a point vector $\mathbf{c}^{i}\in\mathbb{R}^{d}$ in intensional space. Additionally, the designed approach is expected to retain isA transitivity [20] to some degree, namely, InstanceOf-SubClassOf: $(i,r_{e},c_{1})\in T_{e}\wedge(c_{1},r_{c},c_{2})\in T_{c}\to(i,r_{e},c_{2})\in T_{e}$ and SubClassOf-SubClassOf: $(c_{1},r_{c},c_{2})\in T_{c}\wedge(c_{2},r_{c},c_{3})\in T_{c}\to(c_{1},r_{c},c_{3})\in T_{c}$.
@@ 4.1.1 Ontology representation in extensional space
Extensional knowledge provides the information about the types of instances and relationships between instances. A natural idea of modeling extensional knowledge is to use geometric regions to represent concepts, and point vectors to represent instances. For a concept $c$ and its instances ${i_{1},i_{2},\cdots,i_{n}}$, the instance embeddings can be modeled by the vectors $\mathbf{i_{1}},\mathbf{i_{2}},\cdots,\mathbf{i_{n}}$ being inside $c$’s region embedding $\mathbf{G}_{c}$ [20]. This kind of geometry-based method also performs well in capturing SubClassOf relation which belongs to intensional knowledge, but it can’t capture more nuanced and comprehensive information of intensional knowledge. In this work, we follow [43] to represent concept as ellipsoid and instance as point vector in the extensional space, shown in the left part of Figure 3. This representation method considers the anisotropy of concept embedding and has a good capability to represent ontological concepts [43]. For each concept $c\in C$, we represent it as ellipsoid $G_{c}(\mathbf{c},\mathbf{b})$ in extensional space:
@@ 4.1.2 Ontology representation in intensional space
where $\mathbf{c}^{i}\in\mathbb{R}^{d}$, $text_{c}$ denotes the name or concatenation of name and description of concept $c$, $E$ denotes the encoder of the used pretrained language model. In our experiment, we replace the underscores in the concept name with spaces and strip off the brackets on both sides to get the text to be input into the encoder. For instance, concept “<wikicat_Danish_male_film_actors>” will be transformed into ”wikicat Danish male film actors”. We use Sentence-BERT [29], the most popular model for computing text similarity, as the encoder to get the intensional embeddings of ontological concepts. In the experiments, in order to facilitate the comparison, we also conduct experiments of initializing the intensional embeddings of the concepts with random matrix instead of pre-trained model as ablation study. Variants initializing the intensional embeddings directly with random matrix are denoted with UNP, and variants of encoding with pre-trained model are denoted with PRE.
@@ 4.2 Scoring function
In the above formular, $\left\|\mathbf{c}_{\mathbf{i}}^{i}\right\|-\left\|\mathbf{c}_{\mathbf{j}}^{i}\right\|$ can also be used to retain transitivity of concept subsumption in intensional space to some degree [9]. The loss function of SubClassOf Triple can be represented as:
@@
Relational Triple Representation. We define loss function for Relational triples only in extensional space. For a relational triple $(h,r,t)$, we utilize the widely used TransE [2] model to calculate the loss function:
@@ 4.3 Training method
$L_{sub}=\sum_{\delta\in S_{sub}}\sum_{\delta^{\prime}\in S_{sub}^{\prime}}\left[\gamma_{sub}+f_{sub}(\delta)-f_{sub}\left(\delta^{\prime}\right)\right]_{+}$
where $[x]_{+}=\max(0,x)$ and $\gamma_{sub}$ is the margin separating positive triplets and negative triplets. Similarly, for InstanceOf triples and Relational triples, The ranking loss will be:
$L_{ins}$ $=\sum_{\delta\in S_{ins}}\sum_{\delta^{\prime}\in S_{ins}^{\prime}}\left[\gamma_{ins}+f_{ins}(\delta)-f_{ins}\left(\delta^{\prime}\right)\right]_{+}$
@@
The goal of training our proposed model is to minimize the above function, and iteratively update embeddings of concepts, instances, and relations. Every triple in our training set has a label to indicate whether the triple is positive or negative. We follow [2] to generate negative samples.
@@
5 Experiments
5.1 Datasets
We utilized the same datasets, YAGO39K, M-YAGO39K and DB99K-242 in [43]. YAGO39K was constructed by randomly extracting triples from YAGO, encompassing relations such as InstanceOf, SubclassOf, and instance relations, and M-YAGO39K, on the foundation of YAGO39K, was generated by transitivity utilizing the is-A relation to create additional triples [20]. DB99K-242 [11] is created by removing other relations between concepts except the SubClassOf relations on DB [^1] K-174 extracted from DBpedia in [43]. Details of YAGO39K, M-YAGO39K and DB99K-242 are presented in Table 1.
5.2 Experimental Settings
Followed by [43], we selected 13 state-of-the-art knowledge representation learning models as baselines for evaluating our proposed approach. These models are primarily categorized into two classes: (1) Methods distinguishing concepts and instances, including TransC [20], TransFG [44], JECI [35], TransEllipsoid [43] and TransCuboid [43], (2) methods that do not distinguish between concepts and instances, such as TransE [2], TransH [38], TransR [19], TransD [14], DistMult [40], HolE [24], ComplEx [34] and Concept2Vec [28].
The data, codes and determined hyper-parameters can be found at this link [^3] https://github.com/sky-fish23/EIKE.
5.3 Triple Classification
@@
Table 2: Experimental results of InstanceOf triple classification(%). Results of * are taken from [43]

### v7

# Embedding Ontologies via Incoprorating Extensional and Intensional Knowledge

@@ Abstract
Ontologies contain rich knowledge within domain, which can be divided into two categories, namely extensional knowledge and intensional knowledge. Extensional knowledge provides information about the concrete instances that belong to specific concepts in the ontology, while intensional knowledge details inherent properties, characteristics, and semantic associations among concepts. However, existing ontology embedding approaches fail to take both extensional knowledge and intensional knowledge into fine consideration simultaneously. In this paper, we propose a novel ontology embedding approach named EIKE (Extensional and Intensional Knowledge Embedding) by representing ontologies in two spaces, called extensional space and intensional space. EIKE presents a unified framework for embedding instances, concepts and their relations in an ontology, applying a geometry-based method to model extensional knowledge and a pretrained language model to model intensional knowledge, which can capture both structure information and textual information. Experimental results show that EIKE significantly outperforms state-of-the-art methods in three datasets for both triple classification and link prediction, indicating that EIKE provides a more comprehensive and representative perspective of the domain.
@@ 1 Introduction
Ontologies, such as YAGO Suchanek et al. (2007), WordNet Miller (1995) and BioPortal Noy et al. (2009), serve as pivotal tools for organizing, representing, and semantically linking heterogeneous information. They provide a structured framework that defines concepts and their relationships, offering a structured knowledge representation that supports various applications like information retrieval Munir and Anjum (2018) Jun-feng et al. (2005) and semantic web technologies Jonquet et al. (2023)d’Aquin (2023)Wang et al. (2024).
@@
Within an ontology, knowledge can be divided into two categories, intensional knowledge, acting as a terminology or taxonomy, describing general properties, characteristics of concepts and their semantic relationships, and extensional knowledge, also known as assertional knowledge, which pertains to specific individuals within the domain and provides type information about the concrete instances that belong to a specific concept Nardi et al. (2003) d’Aquin (2023). Figure 1 shows a sketch of such an ontology. Comprehensively representing both extensional knowledge and intensional knowledge within an ontology is essential for a nuanced, accurate and adaptable understanding of the domain d’Aquin (2023).
Ontology embedding involves transforming each component in an ontology into continuous and low-dimensional vectors while preserving their semantic relationships. Recently, ontology embedding techniques have been widely investigated and applied to various applications such as semantic search and reasoning Ben Mustapha et al. (2010). Integrating both extensional knowledge and intensional knowledge enriches ontology embedding and provides a more comprehensive and representative perspective of the domain, resulting in embeddings that better capture semantic relationships and similarities among instances and concepts, as well as better apply to applications such as materialization and classification. However, existing ontology embedding models fail to take both of these two types of knowledge into fine consideration simultaneously. For example, existing geometry-based approaches such as Lv et al. (2018)Yu et al. (2023) represent extensional knowledge well, but perform poorly in capturing intensional knowledge since they ignore the inherent characteristics and properties of the concepts. Especially, they can’t capture lexical and textual information. Concept2Vec Qiu and Wang (2020) utilizes a random walk strategy to learn concept embedding through the similarity between concepts, ignoring the influence of a large number of instances. OWL2Vec* Chen et al. (2021) considers the semantic aspects of graph structures, lexical information, and logical constructors, making it suitable for modeling intensional knowledge within an ontology. However, despite the fact that OWL2Vec* takes instances into account, its random walk strategy fail to clearly distinguish the differences between concepts and instances, making it hard to sufficiently model extensional knowledge.
@@ 2.1 Instance-level ontology embedding
Early knowledge embedding methods, called factual knowledge graph embedding, treat both concepts and instances as instances. Mainstream models are mainly divided into translational distance models such as TransE Bordes et al. (2013), TransH Wang et al. (2014), TransR Lin et al. (2015), TransD Ji et al. (2015), and bilinear models, such as RESCAL Nickel et al. (2011), HolE Nickel et al. (2016), DistMult Yang et al. (2014) and ComplEx Trouillon et al. (2016).
Some recent works apply pre-trained language models (PLMs) to knowledge embedding to capture textual information. For example, PretrainKGE Zhang et al. (2020) utilizes BERT Devlin et al. (2018) as encoders to derive embeddings of the entities and relations with descriptions, and input them into structured-based knowledge graph embedding models for training. KEPLER Wang et al. (2021) uses PLMs as encoders to derive description-based embeddings and is trained on the objectives of knowledge embedding (KE) and masked language modeling (MLM). Xie et al. (2022) Chen et al. (2022) both apply sequence-to-sequence framework and pretrained language models to Knowledge Graph Completion (KGC). More recently, Yao et al. (2023) explores Large Language Models for KGC. These methods can leverage descriptive information in knowledge graphs but cannot sufficiently represent concepts and model hierarchical structure.
@@ 2.2 Concept-level ontology embedding
Concept-level ontology embedding methods consider both instances and concepts in an ontological knowledge base as concepts. For instance, EL Embedding Kulmanov et al. (2019) constructs specific scoring functions and loss functions from logical axioms from EL++ by transforming logical relations into geometric relations. Onto2Vec Smaili et al. (2018) and OPA2Vec Smaili et al. (2019) are two ontology embedding algorithms based on word embeddings using skip-gram or CBOW architectures. Concept2Vec Qiu and Wang (2020) utilizes a random walk strategy to learn concept embeddings through the similarity between concepts. CosE Gao et al. (2019) utilizes angle-based and distance-based approaches to design two scoring functions to learn SubClassOf and DisjointWith while preserving transitivity and symmetry. Concept-level ontology embedding focuses on ontological concepts, disregarding instances within the knowledge base, hence its limited scope. Although some ontology embedding methods like OWL2Vec* Chen et al. (2021) take instances into account, they fail to distinguish concepts and instances clearly because of the nature of their learning strategies like random walk. In contrast, our primary focus revolves around methodologies that differentiate concepts and instances.
@@ 2.3 Differentiating concepts and instances for ontology embedding
Recently, some studies have attempted to differentiate between concepts and instances in knowledge representation learning. TransC Lv et al. (2018) represents each concept as a spherical region and each instance as a point vector in the same vector space, capable of handling the isA transitivity. TransFG Yu et al. (2019) adopts TransC’s geometric modeling approach but embeds instances and concepts into different vector spaces to address instance ambiguity. To address the anisotropy issue in TransC and TransFG, Yu et al. (2023) introduces TransEllipsoid and TransCuboid, representing concepts respectively as ellipsoidal and cuboidal regions. Additionally, JOIE Hao et al. (2019) and DGS Iyer et al. (2022) learn embeddings from two-view knowledge graphs, yet they lacks interpretability for the isA transitivity. Zhou et al. (2023) and RMPI Geng et al. (2023) show utilizing ontology-view information can enhance instance-view knowledge representation for applications like knowledge graph completion and inductive relation inference.
Different from previous methods like TransFG Yu et al. (2019) and JOIE Hao et al. (2019) which embed ontologies in two spaces, one space only for concepts and the other only for instances, EIKE embeds extensional knowledge and intensional knowledge in two distinct spaces respectively. This approach represents ontologies more sufficiently since it combine structure-based embeddings and textual-based embeddings. Additionally, EIKE is the first to finely consider two aspects of semantics of a concept: (1) a hight level-summary of its instances; (2) its inherent properties, characteristics and associations with other concepts.
@@ 3 Task Definition
Ontology. Based on the problem formulation in Lv et al. (2018) Yu et al. (2023), we incorporate considerations for literals and descriptions. We consider the relation between concepts (i.e., SubClassOf), the relation between concepts and intances (i.e., InstanceOf), and the relations between instances (for convinience, we call them instance relations), and textual descriptions in ontologies. Ontology is simply formalized as $O=\{C,I,R,T,L\}$. $C$ and $I$ denote the sets of concepts and instances respectively. Relation set $R$ is categorized into three types as $R=\{r_{e},r_{c}\}\cup R_{l}$, where $r_{e}$ and $r_{c}$ denote InstanceOf relation and SubClassOf relation respectively, and $R_{l}$ is the relation set containing instance relations. The triple set $T$ is divided into three disjoint subsets: Relational triple set $T_{l}=\{(i_{m},r_{l},i_{n})\mid i_{m},i_{n}\in I,r_{l}\in R_{l}\}$, SubClassOf triple set $T_{c}=\{(c_{m},r_{c},c_{n})\mid c_{m},c_{n}\in C\}$ and InstanceOf triple set $T_{e}=\{(i,r_{e},c)\mid i\in I,c\in C\}$. $L$ denotes the set of existing descriptions about certain concepts in ontology.
@@
Task Definition. We embed ontologies in extensional space and intensional space. Given an ontology $O$, we map each instance $i$ into a point vector $\mathbf{i}^{e}\in\mathbb{R}^{d}$ and each concept c into a geometric region ${G}(\mathbf{c},\mathbf{b})$ in extensional space and map $\mathbf{i}^{e}$ into a point vector $\mathbf{i}^{i}\in\mathbb{R}^{d}$ and concept $c$ into a point vector $\mathbf{c}^{i}\in\mathbb{R}^{d}$ in intensional space. Additionally, the designed approach is expected to retain isA transitivity Lv et al. (2018) to some degree, namely, InstanceOf-SubClassOf: $(i,r_{e},c_{1})\in T_{e}\wedge(c_{1},r_{c},c_{2})\in T_{c}\to(i,r_{e},c_{2})\in T_{e}$ and SubClassOf-SubClassOf: $(c_{1},r_{c},c_{2})\in T_{c}\wedge(c_{2},r_{c},c_{3})\in T_{c}\to(c_{1},r_{c},c_{3})\in T_{c}$.
@@ 4.1.1 Ontology representation in extensional space
Extensional knowledge provides the information about the types of instances and relationships between instances. A natural idea of modeling extensional knowledge is to use geometric regions to represent concepts, and point vectors to represent instances. For a concept $c$ and its instances $i_{1},i_{2},\cdots,i_{n}$}, the instance embeddings can be modeled by the vectors $\mathbf{i_{1}},\mathbf{i_{2}},\cdots,\mathbf{i_{n}}$ being inside $c$’s region embedding $\mathbf{G}_{c}$ Lv et al. (2018). This kind of geometry-based method also performs well in capturing SubClassOf relation which belongs to intensional knowledge, but it can’t capture more nuanced and comprehensive information of intensional knowledge. In this work, we follow Yu et al. (2023) to represent concept as ellipsoid and instance as point vector in the extensional space, shown in the left part of Figure 3. This representation method considers the anisotropy of concept embedding and has a good capability to represent ontological concepts Yu et al. (2023). For each concept $c\in C$, we represent it as ellipsoid $G_{c}(\mathbf{c},\mathbf{b})$ in extensional space:
@@ 4.1.2 Ontology representation in intensional space
where $\mathbf{c}^{i}\in\mathbb{R}^{d}$, $text_{c}$ denotes the name or concatenation of name and description of concept $c$, $E$ denotes the encoder of the used pretrained language model. In our experiment, we replace the underscores in the concept name with spaces and strip off the brackets on both sides to get the text to be input into the encoder. For instance, concept “<wikicat_Danish_male_film_actors>” will be transformed into ”wikicat Danish male film actors”. We use Sentence-BERT Reimers and Gurevych (2019), the most popular model for computing text similarity, as the encoder to get the intensional embeddings of ontological concepts. In the experiments, in order to facilitate the comparison, we also conduct experiments of initializing the intensional embeddings of the concepts with random matrix instead of pre-trained model as ablation study. Variants initializing the intensional embeddings directly with random matrix are denoted with UNP, and variants of encoding with pre-trained model are denoted with PRE.
@@ 4.2 Scoring function
In the above formular, $\left\|\mathbf{c}_{\mathbf{i}}^{i}\right\|-\left\|\mathbf{c}_{\mathbf{j}}^{i}\right\|$ can also be used to retain transitivity of concept subsumption in intensional space to some degree Gao et al. (2019). The loss function of SubClassOf Triple can be represented as:
@@
Relational Triple Representation. We define loss function for Relational triples only in extensional space. For a relational triple $(h,r,t)$, we utilize the widely used TransE Bordes et al. (2013) model to calculate the loss function:
@@ 4.3 Training method
$L_{sub}=\sum_{\delta\in S_{sub}}\sum_{\delta^{\prime}\in S_{sub}}\left[\gamma_{sub}+f_{sub}(\delta)-f_{sub}\left(\delta^{\prime}\right)\right]_{+}$
where $x]_{+}=\max(0,x)$ and $\gamma_{sub}$ is the margin separating positive triplets and negative triplets. Similarly, for InstanceOf triples and Relational triples, The ranking loss will be:
$L_{ins}$ $=\sum_{\delta\in S_{e}}\sum_{\delta^{\prime}\in S_{e}^{\prime}}\left[\gamma_{ins}+f_{ins}(\delta)-f_{ins}\left(\delta^{\prime}\right)\right]_{+}$
@@
The goal of training our proposed model is to minimize the above function, and iteratively update embeddings of concepts, instances, and relations. Every triple in our training set has a label to indicate whether the triple is positive or negative. We follow Bordes et al. (2013) to generate negative samples.
@@
Table 2: Experimental results of InstanceOf triple classification(%). Results of * are taken from Yu et al. (2023)
5 Experiments
5.1 Datasets
We utilized the same datasets, YAGO39K, M-YAGO39K and DB99K-242 in Yu et al. (2023). YAGO39K was constructed by randomly extracting triples from YAGO, encompassing relations such as InstanceOf, SubclassOf, and instance relations, and M-YAGO39K, on the foundation of YAGO39K, was generated by transitivity utilizing the is-A relation to create additional triples Lv et al. (2018). DB99K-242 Hao et al. (2019) is created by removing other relations between concepts except the SubClassOf relations on DB [^1] K-174 extracted from DBpedia in Yu et al. (2023). Details of YAGO39K, M-YAGO39K and DB99K-242 are presented in Table 1.